# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

### v2

@@ Appendix D Results
Table 17: Summary of memory consumption of datasets used for node-level tasks. All units are in MegaBytes (MB)
@@
$0.1$ $0.798\pm 0.008$ $\mathbf{0.824\pm 0.003}$
$0.3$ $0.798\pm 0.008$ $\mathbf{0.807\pm 0.004}$
$0.5$ $0.807\pm 0.004$ $\mathbf{0.814\pm 0.004}$
$0.7$ $0.797\pm 0.007$ $\mathbf{0.808\pm 0.006}$
Table 18: Average of top 10 accuracies (higher the better) of node classification task on Cora dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the superior performance of the Cluster Nodes method over Extra Nodes.
@@
0.1 $0.504\pm 0.004$ $\mathbf{0.460\pm 0.003}$
0.3 $0.579\pm 0.006$ $\mathbf{0.499\pm 0.006}$
0.5 $0.585\pm 0.004$ $\mathbf{0.505\pm 0.004}$
0.7 $\mathbf{0.504\pm 0.006}$ $0.560\pm 0.004$
Table 19: Average of best 10 losses (lesser the better) of node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the better performance of the Cluster Nodes method over Extra Nodes.
Dataset Coarsening Ratio Top 10 Loss
Chameleon Baseline $0.844\pm 0.000$
$0.1$ $\mathbf{0.460\pm 0.002}$
$0.3$ $0.499\pm 0.006$
$0.5$ $0.587\pm 0.000$
$0.7$ $0.591\pm 0.080$
Crocodile Baseline $0.853\pm 0.000$
$0.1$ $\mathbf{0.362\pm 0.002}$
$0.3$ $0.364\pm 0.001$
$0.5$ $0.375\pm 0.001$
$0.7$ $0.384\pm 0.002$
Squirrel Baseline $0.809\pm 0.000$
$0.1$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.672\pm 0.003$
$0.5$ $0.729\pm 0.004$
$0.7$ $0.694\pm 0.004$
Table 20: Node Regression Results with Gs-train-to-Gs-infer experimental setup.

### v3

@@ Appendix F Results
Table 11: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)
Methods Models Reduction Ratio (r) Dataset
Chameleon Crocodile Squirel
Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
Table 12: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
Methods Models Reduction Ratio (r) Dataset
Cora Citeseer Pubmed DBLP Physics
Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $0.933\pm 0.000$
GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
SGGC GCN 0.1 $0.752\pm 0.002$ $0.704\pm 0.003$ $0.730\pm 0.004$ $0.728\pm 0.019$ $0.892\pm 0.010$
0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
0.7 $0.805\pm 0.002$ $0.711\pm 0.003$ $0.786\pm 0.003$ $0.733\pm 0.014$ $0.931\pm 0.003$
GAT 0.1 $0.630\pm 0.001$ $0.605\pm 0.025$ $89\pm 0.033 0.606$$06\pm 0.017$ $0.854\pm 0.008$
0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
0.7 $0.643\pm 0.019$ $0.570\pm 0.041$ $0.733\pm 0.023$ $0.613\pm 0.018$ $0.851\pm 0.026$
GCOND GCN 0.1 $0.789\pm 0.003$ $0.676\pm 0.020$ $0.785\pm 0.003$ $0.718\pm 0.007$ $0.884\pm 0.008$
0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
0.7 $0.793\pm 0.000$ $0.680\pm 0.012$ $0.775\pm 0.002$ $0.751\pm 0.000$ $0.914\pm 0.004$
GAT 0.1 $0.657\pm 0.067$ $0.553\pm 0.116$ $0.511\pm 0.079$ $0.475\pm 0.057$ OOM
0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
0.7 $0.706\pm 0.062$ $0.453\pm 0.118$ $0.416\pm 0.036$ $0.462\pm 0.033$ OOM
BONSAI GCN 0.1 $0.729\pm 0.008$ $0.576\pm 0.006$ $0.600\pm 0.098$ $0.699\pm 0.031$ $0.822\pm 0.014$
0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
0.7 $0.705\pm 0.009$ $0.582\pm 0.007$ $0.474\pm 0.020$ $0.694\pm 0.023$ $0.844\pm 0.015$
GAT 0.1 $0.685\pm 0.004$ $0.589\pm 0.002$ $0.446\pm 0.014$ $0.642\pm 0.015$ $0.803\pm 0.005$
0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
0.7 $0.703\pm 0.010$ $0.594\pm 0.002$ $0.497\pm 0.028$ $0.705\pm 0.047$ $0.799\pm 0.010$
FIT-GNN GCN 0.1 $0.790\pm 0.004$ $0.701\pm 0.001$ $0.757\pm 0.002$ $0.755\pm 0.001$ $0.909\pm 0.000$
0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $\mathbf{0.932\pm 0.000}$
0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
0.7 $0.813\pm 0.002$ $0.664\pm 0.001$ $0.756\pm 0.002$ $0.746\pm 0.001$ $0.925\pm 0.000$
GAT 0.1 $0.773\pm 0.021$ $0.694\pm 0.003$ $0.737\pm 0.002$ $0.729\pm 0.010$ $0.887\pm 0.003$
0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
0.7 $0.797\pm 0.003$ $0.662\pm 0.003$ $0.759\pm 0.002$ $0.739\pm 0.002$ $0.903\pm 0.003$
Table 13: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
Methods Models Reduction Ratio (r) Dataset
ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $\mathbf{0.524}$
0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
0.5 $0.620$ $\mathbf{0.693}$ $1.227$ $0.475$ $0.561$
0.7 $0.658$ - - - -
SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
0.3 $0.671$ - - - -
0.5 $0.695$ - - - -
0.7 $0.645$ - - - -
GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
0.3 $0.669$ - - - -
0.5 $0.631$ - - - -
0.7 $0.658$ - - - -
Table 14: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Table 15: The Tables show the comparison between Extra Nodes and Cluster Nodes methods for node classification and node regression task. We show the average of the top 10 accuracies (higher the better) of the node classification task on Cora dataset and the average of the best 10 losses (lesser the better) of the node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup. They demonstrate the superior performance of the Cluster Nodes method over Extra Nodes.
(a) Node Classification on Cora.
@@
$0.1$ $\mathbf{0.798\pm 0.008}$ $0.790\pm 0.004$
$0.3$ $0.798\pm 0.008$ $\mathbf{0.800\pm 0.003}$
$0.5$ $0.807\pm 0.004$ $\mathbf{0.829\pm 0.002}$
$0.7$ $0.797\pm 0.007$ $\mathbf{0.813\pm 0.002}$
(b) Node Regression on Chameleon.
@@
0.1 $\mathbf{0.504\pm 0.004}$ $0.506\pm 0.006$
0.3 $0.579\pm 0.006$ $\mathbf{0.489\pm 0.003}$
0.5 $0.585\pm 0.004$ $\mathbf{0.484\pm 0.004}$
0.7 $\mathbf{0.504\pm 0.006}$ $0.539\pm 0.009$
Table 16: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
PROTEINS ZINC (subset)
variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.573}$ $\mathbf{0.620}$
algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
Cora Chameleon
variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$